OpenStudy (anonymous):

Serenity, Lucy, and Piper hiked for 4 hours. They hiked 1\frac{3}{4}miles each hour for the first three hours. In the fourth hour, they hiked \frac{3}{10}of the total distance that they hiked during the first three hours. What was the total distance the girls \frac{31}{40} \frac{5}{6} \frac{5}{6} \frac{33}{40}

OpenStudy (just_one_last_goodbye):

Hello @Springtrap550 I can guarantee you that @texaschic101 's help will be very helpful ^_^ I dont know @Kreayhea_ but i'm sure she will get my recommendation depending on her performance

OpenStudy (anonymous):

1.25x3=5.25 5.25 divided by .3 = 17.5 >(Convert into a fraction and you should get an answer, if not apologies.)

OpenStudy (anonymous):

^ 1.75*

OpenStudy (texaschic101):

1 3/4 miles for 3 hrs... 7/4 * 3 = 21/4 On the 4th mile...they hiked 3/10 of total distance.. 3/10 of 21/4 = 3/10 * 21/4= 63/40 I am doing something wrong...my answer is not there

OpenStudy (texaschic101):

oh wait....total distance... 21/4 + 63/40 = 210/40 + 63/40 = 273/40 = 6 33/40 total miles